OpenStudy (anonymous):

Please click to see equation. Medal will be rewarded to most helpful. Please explain!

OpenStudy (anonymous):



OpenStudy (austinl):

\[\frac{x}{6} = -21\] \[(6)\times \frac{x}{6} = -21\times (6)\] \[x = -21\times (6)\]

OpenStudy (anonymous):

Thank you very much! :)

OpenStudy (austinl):

No problemo!